OpenStudy (anonymous):

find the domain of the function algebraicly. : f(x)= square root (x^2+4) . can someone please explain to me how to do this ?

OpenStudy (mathmate):

The domain (i.e. possible values of x) of a polynomial is all real, i.e. (-inf,+inf). However, the square-root function sometimes puts a limit to what values x can take, in such a way that the expression within the square-root radical must be positive. In the given problem, the expression in the sqrt radical is positive for all values of x, so the domain for f(x) is (-inf,+inf), or all real.

OpenStudy (anonymous):

thank you so much ! i actually understand this now lol

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

what about f(X)= 3x-1/(x-3)(x-1)

OpenStudy (mathmate):

Cannot answer until I know what f(x) looks like. Sometimes parentheses are missed out which makes a different function! The above function as is is the same as: \[f(x)=\frac{(3x-1)(x-1)}{(x-3)}\] But if missing parentheses are put around both numerator and denominator, f(X)= (3x-1)/((x-3)(x-1))it means: \[f(x) = \frac{(3x-1)}{(x-3)(x-1)}\]

OpenStudy (anonymous):

oh its the second one

OpenStudy (mathmate):

In any case, this is similar to the square-root case, except that here what we don't want to see is that the denominator becomes zero. In the first interpretation, the rational function has a domain of all real, except when the denominator becomes zero, which happens when x=3. So the domain is (-inf,3)U(3,+inf). This interval notation excludes 3 from the domain. "(" means exclude the value, while [ or ] includes the value. Sometime this is also written as (-inf,+inf)\{3} which means all real excluding 3. The answer to the second interpretation can be found in a similar way.

OpenStudy (anonymous):

how would you do it algebraically ?

OpenStudy (mathmate):

You can actually read it off the second interpretation by considering when the denominator will be zero.

OpenStudy (mathmate):

It would be all real except for two values which make the denominator vanish.

OpenStudy (anonymous):

ohh thanks your the bestt

OpenStudy (mathmate):

You're welcome! :)